OpenStudy (anonymous):

I am trying to find help in dividing polynomials when there are two variables : example problem (4x^3y^2+8xy^2-12x^2y^3) / 4xy

OpenStudy (anonymous):

x^2y + 2y -3xy^2

OpenStudy (anonymous):

How?

OpenStudy (gw2011):

Divide each term by 4xy. This is how you can do it: 4x^3y^2/4xy = x^2y 8xy^2/4xy = 2y -12x^2y^3/4xy = -3xy^2 Therefore, the answer is: x^2y+2y-3xy^2